You are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.

  • Consumption 350 billion G
  • Transfer payments 100 billion G
  • Investment 100 billion G
  • Government purchases 200 billion G
  • Exports 50 billion G
  • Imports 150 billion G
  • Bond purchases 200 billion G
  • Earnings on foreign investments 75 billion G
  • Foreign earnings on Amagre investment 25 billion G
  1. Compute net foreign investment.
  2. Compute net exports.
  3. Compute GDP.
  4. Compute GNP.

In addition to responding with a quantitative answer, briefly describe how you arrived at your answers.


1
Expert's answer
2021-12-16T16:19:08-0500

1.Net investment is gross investment minus depreciation

In=Ig-A

In=100 billion G

2.Net exports is the difference between a country's exports and imports

Xn=Ex-Im=50-100=-150

3.Gross domestic product (GDP) is the sum of the values of all goods and services produced in the state.

"GDP = C + I + G + Xn=350+100+200-150=500"

4.

"GNP=GNP+\Delta=500+75+25-200=400"

where Δ = (primary income received by residents abroad) — (primary income received by non-residents in the economic territory of the country); that is, the balance of the exchange of primary income with the rest of the world.